# Evolving Generalizable Parallel Algorithm Portfolios via Domain-Agnostic Instance Generation

### v3

@@ V-C Experimental Protocol
Following the experimental protocol in the CEPS paper [11], the number of member algorithms in PAP, i.e., $K$, was set to $4$, and solution quality was used as the performance indicator $f$. For both CEPS and DACE, identical parameter settings were used to make fair comparisons. Specifically, the number of co-evolution rounds (i.e., $MaxRound$) was set to 4, configuration mining was repeated 20 times (i.e., $n=20$), and the maximum number of trials in SMAC was set to 1600. Both methods used the same randomly sampled initial configuration set $C$ in their initialization phase. For instance generation in both methods, the mutation operator was run for a maximum of 200 iterations, and the hardest instance (on which the PAP achieves the lowest solution quality) from these iterations was returned to update the instance population. The hyper-parameters of the NIR structure in DACE are shown in Table I. The weighting hyper-pamraeters $\lambda_{1}$ and $\lambda_{2}$ in the NIR training loss function were set to 1 and 0.0005, respectively.
@@ V-D Test Results and Analysis
Given that the key distinction between CEPS and DACE lies in their instance generators, we speculate that the superior performance of DACE’s domain-agnostic instance mutation operator might be attributed to two aspects. First, by transforming instance generation into a continuous optimization problem through NIRs as described in Alg. 1, DACE can leverage powerful continuous optimization methods to identify challenging instances more effectively than CEPS’s domain-specific approaches. This advantage is particularly evident in CAOP, where CEPS’s instance mutation operator failed to identify any instances more challenging than the initial training instances, resulting in no expansion of the training instance population. In contrast, DACE’s NIR-based mutation operator successfully identified new challenging instances that enriched the problem instance population. Second, NIR-based representations could lead to the generation of instances as diverse as those generated by CEPS’s domain-specific generators, contributing to effective PAP generalization. This comparable diversity is also validated in Section V-F. These results demonstrate that DACE not only successfully eliminates the need for domain-specific instance generators but also achieves superior effectiveness.
@@
Another observation from Fig. 4 is that DACE and CEPS, which incorporate instance generation mechanisms, consistently outperform GLOBAL and PARHYDRA, which lack such mechanisms. This indicates generating synthetic instances during PAP construction effectively improves generalization in few-shot scenarios, aligning with findings from previous work [11]. Moreover, GLOBAL and PARHYDRA perform worse than the manually constructed BRKGA-PAP across most problem classes. This performance degradation is likely due to overtuning, where the PAPs become overly specialized to the limited training set, compromising their generalization capabilities.
@@ V-E Analysis of PAP’s Generalization through Co-Evolution
An interesting observation spans all three problem classes. While PAPs were constructed using training instances of only certain dimensions (30 for CCP, 80 for CIMP and CAOP), they were tested on both matching dimensions and larger ones (40 for CCP, 100 for CIMP and CAOP). DACE’s performance improvement is particularly noticeable on higher-dimensional test instances in CAOP and CCP, while for CIMP the improvement is consistent across both dimensions. This observation has two important implications. First, it suggests that higher-dimensional instances, which are typically more challenging to solve with limited solution evaluations, provide more opportunities for performance improvement. Second, it demonstrates DACE’s ability to construct PAPs that can effectively handle problems of varying dimensions, even when trained only on instances of fixed dimensions.
@@ V-F Visual Analysis of NIR-based Instance Generation
To address RQ3 raised at the beginning of this section, experiments were conducted to examine whether the generated NIRs can effectively represent problem instances in the problem class. A visualization method was developed to compare the distributions of instances from different sources across three problem classes. The visualization included instances from training sets, test sets, and those generated by DACE and CEPS. Additionally, to investigate the role of domain-invariant features in NIRs, a variant of DACE called $\text{DACE}_{NoReg}$ was introduced. Unlike DACE which mutates instance embeddings and then uses a hypernetwork to generate scorer weights, $\text{DACE}_{NoReg}$ directly applies the mutation operator from Alg. 1 to modify the scorer weights. This removal of the hypernetwork, a key component for capturing domain-invariant features, allows examination of its importance in instance generation.
@@
A visualization method was developed to project problem instances into a 2D space based on features extracted from their solution-to-objective value mappings, with dimensionality reduction performed using t-SNE[47]. The method builds on findings from previous research showing that neighborhood characteristics significantly influence the difficulty of combinatorial optimization problems [48, 49]. Since these characteristics are determined by the objective functions, which varies across problem classes, they serve as effective features for visualizing and distinguishing different problem instances. The feature extraction process began by randomly sampling 1M solutions for each instance and evaluating their normalized objective values. From these solutions, 10M pairs were randomly sampled to analyze the relationship between Hamming distances and objective value differences. The solution pairs were grouped into $m$ sets based on their Hamming distances, with each set containing pairs of equal distance. Using 16 quantiles $\left[a_{0},a_{1},\cdots,a_{15}\right]$ where $a_{i}=\frac{i}{15}$, the set $V_{i}$ was defined as containing pairs with the $\lfloor a_{i}\times m\rfloor$-th largest Hamming distance. For each $V_{i}$, two statistics were calculated: $b_{i}$ as the mean objective value difference and $c_{i}$ as the standard deviation of objective value differences. These statistics were combined into two vectors: $\boldsymbol{b}=\left[b_{0},b_{1},\cdots,b_{15}\right]$ and $\boldsymbol{c}=\left[c_{0},c_{1},\cdots,c_{15}\right]$. The final feature vector for each problem instance is $\boldsymbol{b}\oplus\boldsymbol{c}$, where $\oplus$ is the concatenation operator. After obtaining feature vectors for all instances, t-SNE was applied to reduce their dimensionality to 2. Fig. 6 shows the visualization results of all problem instances used in the experiment, including those from the training set, test set, instances generated by CEPS, as well as NIRs generated by DACE and $\text{DACE}_{NoReg}$.
@@
In the CIMP class (Fig. 6b), all three methods – DACE, CEPS, and $\text{DACE}_{NoReg}$ – generate instances that overlap with the reference area, though their distributions differ. The diversity of instances generated by DACE and CEPS appears comparable, contributing to effective PAP generalization. For the CAOP class (Fig. 6c), only DACE and $\text{DACE}_{NoReg}$ are compared since CEPS failed to generate challenging instances in this class. Specifically, it could not generate instances where the PAP performs worse than on the initial training instances. In Fig. 6c, both DACE and $\text{DACE}_{NoReg}$ show limited overlap with the reference area. Actually, instances generated by $\text{DACE}_{NoReg}$ deviate significantly from the reference area and even overlap with the CCP region, which explains the previously observed overlap between CAOP and CIMP instances. It can be also observed that DACE’s coverage area lies closer to the reference area than $\text{DACE}_{NoReg}$, indicating that problem class regularization helps generate more realistic instances. In the CCP class (Fig. 6d), both DACE and CEPS achieve coverage areas that align with the reference area, though DACE’s alignment and coverage are slightly inferior to CEPS’s. On the other hand, $\text{DACE}_{NoReg}$’s coverage area largely falls outside the reference area, with some instances showing significant deviation. This suggests that without problem class regularization, $\text{DACE}_{NoReg}$ generates instances with substantially different neighborhood characteristics from authentic CCP instances, potentially reducing their usefulness in PAP construction.
The above results demonstrate that generated NIRs in DACE effectively resemble authentic instances in the problem class and exhibit diversity comparable to CEPS, providing a positive answer to RQ3.
@@ VI Conclusion and Discussion
Several promising directions for future research are outlined. First, the effectiveness of NIR-based instance generation could be further improved, particularly for problems like CAOP where the current approach shows limited overlap with actual instances. Second, DACE currently constructs PAPs based on general-purpose EAs like BRKGA, which typically achieve lower search efficiency compared to specialized algorithms that can leverage domain knowledge to guide the search. One promising direction is to integrate transfer optimization mechanism [50] into DACE that enables member algorithms to learn and transfer solution patterns across different problem instances, potentially bridging the efficiency gap between general-purpose and specialized algorithms while maintaining DACE’s domain-agnostic advantage.
@@ 4 $m^{\prime},f^{\prime}\leftarrow m,f\left(P,m\right)$;
6for $iter\leftarrow 1$ to $MaxIter$  do
7       $\boldsymbol{e}_{1},\boldsymbol{e}_{2},\cdots,\boldsymbol{e}_{N}\leftarrow$ sampling $N$ weights from $\mathcal{N}\sim\left(\boldsymbol{\mu},\boldsymbol{\sigma}\right)$ randomly;
@@ 20       $\boldsymbol{f}^{M}\leftarrow\left[f_{1}-f_{N+1},f_{2}-f_{N+2},\cdots,f_{N}-f_{2N}\right]$;
23       $\boldsymbol{f}^{S}\leftarrow\left[\frac{f_{1}+f_{N+1}}{2}-b,\frac{f_{2}+f_{N+2}}{2}-b,\cdots,\frac{f_{N}+f_{2N}}{2}-b\right]$;
@@
In the NIR-based instance mutation operator described in Alg. 1, PGPE [51] is used as the optimizer. The details are shown in Alg. 3. Specifically, PGPE employs the symmetric sampling exploration strategy (lines 6-9) and the strategy update method (lines 16-20). It uses an iteratively updated multivariate Gaussian distribution to explore the vector space of problem instance embeddings. The problem instance embedding vector that has the lowest $f$ value in this exploration process replaces the problem instance embedding vector in $m$, yielding a new NIR $m^{\prime}$ as the newly generated problem instance (line 15). The hyper-parameters in Alg. 3 are set to $\boldsymbol{\sigma}^{init}=\boldsymbol{1}$, $\alpha_{\mu}=0.05$, $\alpha_{\sigma}=0.1$, and $\sigma^{limit}=0.01$. Compared to the recommended configuration [51], we choose a larger $\boldsymbol{\sigma}$ and a lower $\alpha_{\sigma}$ to encourage the operator to find more diverse solutions.
@@ Appendix B Value Ranges of BRKGA’s Parameters
BRKGA [38] is used as the parameterized optimization algorithm in DACE, as mentioned in Section IV. The descriptions and value ranges of BRKGA’s parameters are listed in Table V and Table VI, respectively.
@@
Appendix C BRKGA-PAP and SMARTEST-PAP

### v4

@@ V-C Experimental Protocol
Following the experimental protocol in the CEPS paper [tang_few-shots_2021], the number of member algorithms in PAP, i.e., $K$, was set to $4$, and solution quality was used as the performance indicator $f$. For both CEPS and DACE, identical parameter settings were used to make fair comparisons. Specifically, the number of co-evolution rounds (i.e., $MaxRound$) was set to 4, configuration mining was repeated 20 times (i.e., $n=20$), and the maximum number of trials in SMAC was set to 1600. Both methods used the same randomly sampled initial configuration set $C$ in their initialization phase. For instance generation in both methods, the mutation operator was run for a maximum of 200 iterations, and the hardest instance (on which the PAP achieves the lowest solution quality) from these iterations was returned to update the instance population. The hyper-parameters of the NIR structure in DACE are shown in Table I and the parameters were chosen based on empirical experience. The weighting hyper-pamraeters $\lambda_{1}$ and $\lambda_{2}$ in the NIR training loss function were set to 1 and 0.0005, respectively.
@@ V-D Test Results and Analysis
Given that the key distinction between CEPS and DACE lies in their instance generators, we speculate that the superior performance of DACE’s domain-agnostic instance mutation operator might be attributed to two aspects. First, by transforming instance generation into a continuous optimization problem through NIRs as described in Alg. 1, DACE can leverage powerful continuous optimization methods to identify challenging instances more effectively than CEPS’s domain-specific approaches. In the co-evolutionary training process, an instance $m^{\prime}$ is considered challenging when the solution quality achieved by the PAP $P$ on that instance is inferior to the solution quality achieved by $P$ on any instance in the current population $M$:
$f\left(P,m^{\prime}\right)<\min_{m\in M}f\left(P,m\right),$ (15)
where the solution quality $f$ is defined in Eq. (11). This advantage is particularly evident in CAOP, where CEPS’s instance mutation operator failed to identify any challenging instances, resulting in the break of the instance mining process (lines 24-25 of Algorithm 2) and no expansion of the training instance population. In contrast, DACE’s NIR-based mutation operator successfully identified new challenging instances that enriched the problem instance population. These results demonstrate that DACE not only successfully eliminates the need for domain-specific instance generators but also achieves superior effectiveness.
@@
Another observation from Fig. 4 is that DACE and CEPS, which incorporate instance generation mechanisms, consistently outperform GLOBAL and PARHYDRA, which lack such mechanisms. This indicates generating synthetic instances during PAP construction effectively improves generalization in few-shot scenarios, aligning with findings from previous work [tang_few-shots_2021]. Moreover, GLOBAL and PARHYDRA perform worse than the manually constructed BRKGA-PAP across most problem classes. This performance degradation is likely due to overtuning, where the PAPs become overly specialized to the limited training set, compromising their generalization capabilities.
@@ V-E Analysis of PAP’s Generalization through Co-Evolution
An interesting observation spans all three problem classes. While PAPs were constructed using training instances of only certain dimensions (30 for CCP, 80 for CIMP and CAOP), they were tested on both matching dimensions and larger ones (40 for CCP, 100 for CIMP and CAOP). DACE’s performance improvement is particularly noticeable on higher-dimensional test instances in CAOP and CCP, while for CIMP the improvement is consistent across both dimensions. This observation has two important implications. First, it suggests that higher-dimensional instances, which are typically more challenging to solve with limited solution evaluations, provide more opportunities for performance improvement. Second, it demonstrates that DACE can construct effective PAPs that generalize well to instances of different dimensions, even when trained on a small set of fixed-dimension instances.
@@ V-F Visual Analysis of NIR-based Instance Generation
To address RQ3 raised at the beginning of this section, experiments were conducted to examine whether the generated NIRs can effectively represent problem instances in the problem class. A visualization method was developed to compare the distributions of instances from different sources across three problem classes. The visualization included instances from training sets, test sets, and those generated by DACE and CEPS. Additionally, to investigate the role of domain-invariant features in NIRs, a variant of DACE called $\text{DACE}_{NoReg}$ was introduced. Unlike DACE which mutates instance embeddings and then uses a hypernetwork to generate scorer weights, $\text{DACE}_{NoReg}$ directly applies the mutation operator from Alg. 1 to modify the scorer weights. This removal of the hypernetwork, a key component for capturing domain-invariant features, allows examination of its importance in instance generation. It should be noted that the instances in the training set and test set, as well as the instances generated by CEPS, are indeed authentic problem instances belonging to their specific problem classes. In contrast, the instances generated by DACE and $\text{DACE}_{NoReg}$ are represented as NIRs.
@@
A visualization method was developed to project problem instances into a 2D space based on features extracted from their solution-to-objective value mappings, with dimensionality reduction performed using PCA. The method was inspired by the previous fitness-distance correlation (FDC) research [jones_fitness_1995, altenberg_fitness_1997, pitzer_comprehensive_2012], which demonstrated the significant influence of neighborhood characteristics on the difficulty of combinatorial optimization problems. However, standard FDC focuses on analyzing the correlation between solution fitness and distance specifically to the global optimum. While this correlation is useful for algorithm compatibility analysis, it lacks the information required for our goal of fine-grained problem characterization and classification. Therefore, we propose a novel feature extraction method that captures neighborhood characteristics across the entire global solution space, rather than concentrating solely on the optimum. By statistically characterizing fitness variations relative to distance throughout the solution space, this approach provides richer information. The feature extraction process began by randomly sampling 1M solutions for each instance and evaluating their normalized objective values. From these solutions, 10M pairs were randomly sampled to analyze the relationship between Hamming distances and objective value differences. The solution pairs were grouped into $m$ sets based on their Hamming distances, with each set containing pairs of equal distance. Using 16 quantiles $\left[a_{0},a_{1},\cdots,a_{15}\right]$ where $a_{i}=\frac{i}{15}$, the set $V_{i}$ was defined as containing pairs with the $\lfloor a_{i}\times m\rfloor$-th largest Hamming distance. For each $V_{i}$, two statistics were calculated: $b_{i}$ as the mean objective value difference and $c_{i}$ as the standard deviation of objective value differences. These statistics were combined into two vectors: $\boldsymbol{b}=\left[b_{0},b_{1},\cdots,b_{15}\right]$ and $\boldsymbol{c}=\left[c_{0},c_{1},\cdots,c_{15}\right]$. The final feature vector for each problem instance is $\boldsymbol{b}\oplus\boldsymbol{c}$, where $\oplus$ is the concatenation operator. After obtaining feature vectors for all instances, we construct a dimensionality reduction mapping from 32 to 2 dimensions using PCA, based on only training and test sets. Subsequently, this mapping is applied to reduce all feature vectors to their 2-dimensional representations. Fig. 6 shows the visualization results of all problem instances used in the experiment, including those from the training set, test set, instances generated by CEPS, as well as NIRs generated by DACE and $\text{DACE}_{NoReg}$.
@@
In the CIMP class (Fig. 6b), all three methods – DACE, CEPS, and $\text{DACE}_{NoReg}$ – generate instances that overlap with the reference area, though their distributions differ. For the CAOP class (Fig. 6c), only DACE and $\text{DACE}_{NoReg}$ are compared since CEPS failed to generate challenging instances in this class. Specifically, CEPS cannot generate instances with lower quality than the initial training instances in CAOP, while quality is defined in Eq. (11). In Fig. 6c, both DACE and $\text{DACE}_{NoReg}$ show limited overlap with the reference area. Actually, instances generated by $\text{DACE}_{NoReg}$ deviate significantly from the reference area and even overlap with the CCP region, which explains the previously observed overlap between CAOP and CIMP instances. It can be also observed that DACE’s coverage area lies closer to the reference area than $\text{DACE}_{NoReg}$, indicating that problem class regularization helps generate more realistic instances. In the CCP class (Fig. 6d), both DACE and CEPS achieve coverage areas that align with the reference area, though DACE’s alignment and coverage are slightly inferior to CEPS’s. On the other hand, $\text{DACE}_{NoReg}$’s coverage area largely falls outside the reference area, with some instances showing significant deviation. This suggests that without problem class regularization, $\text{DACE}_{NoReg}$ generates instances with substantially different neighborhood characteristics from authentic CCP instances, potentially reducing their usefulness in PAP construction.
The above results demonstrate that generated NIRs in DACE effectively resemble authentic instances in the problem class, providing a positive answer to RQ3. It is worth noting that the above conclusion’s scope could be limited, since instance similarity was measured only using the selected fitness-distance features. A more comprehensive characterization of instance properties is an important direction for future work. This would involve incorporating other key properties, such as the “challenge” (or “difficulty level”) that the generated instances pose for PAPs.
@@ VI Conclusion and Discussion
When tackling real-world problems, PAPs offer a significant advantage over using a single algorithm, i.e., they can achieve superior solution quality. Naturally, PAPs introduce some additional computational overhead. However, with the widespread availability of modern parallel computing platforms, PAPs provide a straightforward way to leverage these resources, effectively mitigating this overhead. The enhanced solution quality is particularly valuable in quality-sensitive scenarios, such as the CCP problem class explored in this work, where even small improvements can have substantial practical implications.
Several promising directions for future research are outlined. First, the training of NIRs is currently limited to problem instances of a single dimension. While we introduced a simple workaround in the supplementary (Appendix E) to potentially handle small dimensional variations, developing a truly robust cross-dimensional training approach remains a critical and open challenge for future work. This would significantly expand the application scenarios of our method. Second, a more comprehensive characterization of instance properties would be beneficial for analyzing the generated instances (NIRs). One way is to move beyond the current focus on instance similarity to also include a deeper assessment of the “challenge” (or “difficulty level”) that the generated instances pose for PAPs. Such an assessment could be achieved by analyzing the instances’ fitness landscapes or the specific behaviors of solvers when applied to them. Finally, since DACE can be applied to various optimizers beyond BRKGA, extending it to other optimizers, such as the estimation of distribution algorithms (EDA) variants, or even constructing a PAP containing multiple types of optimizers, is a valuable future direction.
Acknowledgments
This work was supported in part by the Guangdong Major Project of Basic and Applied Basic Research under Grant 2023B0303000010, in part by the Natural Science Foundation of China under Grant T2495254, and in part by the National Natural Science Foundation of China under Grant 62502192.
@@ 4 $m^{\prime},f^{\prime}\leftarrow m,f\left(P,m\right)$;
6for $iter\leftarrow 1$ to MaxIterMaxIter do
7  $\boldsymbol{e}_{1},\boldsymbol{e}_{2},\cdots,\boldsymbol{e}_{N}\leftarrow$ sampling NN weights from $\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\sigma}\right)$ randomly;
@@ 20  $\boldsymbol{f}^{M}\leftarrow\left[f_{1}-f_{N+1},f_{2}-f_{N+2},\cdots,f_{N}-f_{2N}\right]$;
23
$\mathbf{f}^{S}\leftarrow\left[\frac{f_{1}+f_{N+1}}{2}-f_{b},\frac{f_{2}+f_{N+2}}{2}-f_{b},\cdots,\frac{f_{N}+f_{2N}}{2}-f_{b}\right]$
@@
In the NIR-based instance mutation operator described in Alg. 1, PGPE [sehnke_parameter-exploring_2010] is used as the optimizer. The details are shown in Alg. 3. Specifically, PGPE employs the symmetric sampling exploration strategy (lines 6-9) and the strategy update method (lines 16-20). It uses an iteratively updated multivariate Gaussian distribution to explore the vector space of problem instance embeddings. The problem instance embedding vector that has the lowest $f$ value in this exploration process replaces the problem instance embedding vector in $m$, yielding a new NIR $m^{\prime}$ as the newly generated problem instance (line 15). The hyper-parameters in Alg. 3 are set to $\boldsymbol{\sigma}^{init}=\boldsymbol{1}$, $\alpha_{\mu}=0.05$, $\alpha_{\sigma}=0.1$, and $\sigma^{limit}=0.01$. Compared to the recommended configuration [sehnke_parameter-exploring_2010], we choose a larger $\boldsymbol{\sigma}$ and a lower $\alpha_{\sigma}$ to encourage the operator to find more diverse solutions.
@@ Appendix B Value Ranges of BRKGA’s Parameters
BRKGA [goncalves_biased_2011] is used as the parameterized optimization algorithm in DACE, as mentioned in Section IV. The descriptions and value ranges of BRKGA’s parameters are listed in Table V and Table VI, respectively.
@@
Appendix C BRKGA-PAP, SMARTEST-PAP and PAPs Constructed Automatically